OpenStudy (anonymous):

Write in slope-intercept form: (-2,3) and (0,1)

OpenStudy (anonymous):

First you have to find the slope of the two points using y2 - y1/x2 - x1 --> 1 - 3 = -2 over 0 - (-2) = -2/2 which equals -1 So now you can put it in slope-intercept form: y = mx + b --> y = -1 + 1 The y-intercept (or the variable b) comes from when the x-value equals zero! (0,1) Is this what you were looking for?

OpenStudy (anonymous):

Yep, so its y=-1x+1?

OpenStudy (anonymous):

Sorry! Forgot the x! But, yup, that should be it :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

No problem at all!